A test for displaying Japanese fonts in dvi by xdvik-22.84.10-0m

平成 18 年 4 月 2 日

1 test

Can you see all entries in the following tabular? If thing goes well, you can see two Japanese sentences in a right colum.

English  Japanese
Serif  This is 1 pen. これは 1 本のペンです。
Sans Serif This is 1 pen これは 1 本のペンです。

I confirmed that the Japanese fonts in dvi file are displayed with X11R6.9.0 which is said to have the same code base as X11R7.0 (by changing %x11dir in xdvik.spec in xdvik-22.84.10-0m to /usr/X11R6/lib/X11 from /usr/share/X11).